Sophie's favorite number is a two-digit number. If she reverses the digits, the result is $45$ less than her favorite number. The sum of the digits in her favorite number is $6$. What is Sophie's favorite number?

Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3817)


For the original number we have
t = tens digit
u = units digit
That constructs 10t+u

For instance if the number is 23 then t = 2 and u = 3.
10t+u = 10*2+3 = 23.

The digits of Sophie's mystery number add to 6.
t+u = 6 which solves to u = 6-t
This will be useful in a substitution later.

10u+t is the result of reversing the digits. This is 45 less than her favorite number.

swapped = original - 45
10u + t = 10t+u - 45
10(u) + t = 10t+u - 45
10(6-t) + t = 10t+6-t - 45
60-10t+t = 10t+6-t-45
60-9t = 9t-39
-9t-9t = -39-60
-18t = -99
t = -99/(-18)
t = 5.5
We do not get a whole number result for the tens digit, so there must be a typo somewhere in your question.
Perhaps the digits add to something else other than 6?
Or maybe the "45" should be another value?
I would ask your teacher for clarification.

Answer by ikleyn(52787)

.
Sophie's favorite number is a two-digit number. If she reverses the digits, the result is $45$ less
than her favorite number. The sum of the digits in her favorite number is $6$. What is Sophie's favorite number?
~~~~~~~~~~~~~~~~~~~~~~~~~


The given info that the difference of the number and the reversed number is  45
tells us that the difference of the digits is  45/9 = 5.

Hence,  the digits are of different parity.

Then the sum of the digits can not be  6.   Contradiction.


        Beware of sending wrong tasks to this forum.

        Avoid sending wrong tasks to this forum.

        Big brother is watching everything you do here.